# PROPHET: An Inferable Future Forecasting Benchmark with Causal Intervened Likelihood Estimation

### v1

@@ 4.2 Causal Intervened Likelihood
For any two occurrences of news, the one that occurs later in date cannot have an effect on the one earlier:
$\forall~{}i,j,$ $if~{}~{}{\mathcal{T}}_{i}\prec{\mathcal{T}}_{j},$ (4)
$then~{}~{}P({\mathcal{X}}_{i}|{\mathcal{X}}_{j})=P({\mathcal{X}}_{i}).$
Second, causal relationships between events that are widely separated in time should be mediated by events that occur between them. We group all the news in chronological order, with a group size representing 10 days passing. $\mathrm{G}({\mathcal{X}}_{i})$ stands for the index of the group in which ${\mathcal{X}}_{i}$ is in. In our case, if $\mathrm{G}({\mathcal{X}}_{i})<\mathrm{G}({\mathcal{X}}_{j})$ indicates ${\mathcal{T}}_{i}\prec{\mathcal{T}}_{j}$, namely the $i^{th}$ news happens before the $j^{th}$.
Assumption 2. $w$-window Dependency
Variables in the $i^{th}$ group can only be directly influenced by variables within the previous $w$ groups (i.e., groups $i-1,i-2,\dots,i-w$). Consequently, there exist no direct edges between ${\mathcal{X}}_{i}$ and ${\mathcal{X}}_{j}$ for any $j$ outside this window:
$\forall~{}i,j,$ $if~{}~{}\mathrm{G}({\mathcal{X}}_{j})-\mathrm{G}({\mathcal{X}}_{i})>w,$ (5)
$then~{}~{}({\mathcal{X}}_{i},{\mathcal{X}}_{j})\notin\text{edges~{}of~{}SCM}.$
Lastly, news in the same group should have no causal relation in between.
Assumption 3. Concurrent Independency
Any two pieces of news that occurred in the same group are independent:
$\forall~{}i,j,$ $if~{}~{}\mathrm{G}({\mathcal{X}}_{j})=\mathrm{G}({\mathcal{X}}_{i}),$ (6)
$then~{}~{}({\mathcal{X}}_{i},{\mathcal{X}}_{j})\notin\text{edges~{}of~{}SCM}.$
With these assumptions, we can derive CIL estimation. We show the calculation of $P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|do({\mathcal{X}}_{i}=1))$, then $P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|do({\mathcal{X}}_{i}=0))$ can be computed similarly.
@@
The intervened probability $P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|do({\mathcal{X}}_{i}=1))$ can be convert into observation probability:
$P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|do({\mathcal{X}}_{i}=1))\doteq P_{m}({\mathcal{Y}}|{\mathcal{X}}_{i}=1)$ (7)
$=$ $\sum_{n_{1},}\cdots P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots)P({\mathcal{X}}_{n_{1}},\cdots)$
$0<\mathrm{G}({\mathcal{X}}_{i})-\mathrm{G}({\mathcal{X}}_{n_{j}})\leqq w,\forall n_{j}.$
# News # Token Max TS Mean TS
100 853.95 31 16.54
Table 1: Statistics of the grounding news. TS stands for time span between the oldest and latest news of a question. The unit is a month.
Models Retrieval Reasoning L1 L2
Brier Score $\downarrow$ CIL $\uparrow$ Brier Score $\downarrow$ CIL $\uparrow$
GPT-4o w.o. RAG ScratchPAD 25.42 $\pm$ 0.09 - 23.09 $\pm$ 1.38 -
Naive RAG 21.22 $\pm$ 0.30  (+4.20) 0.07 $\pm$ 0.00 22.79 $\pm$ 0.64  (+0.30) -4.60 $\pm$ 0.00
APP 20.02 $\pm$ 0.26  (+5.40) 1.47 $\pm$ 0.16 24.25 $\pm$ 0.69  (-1.16) -4.68 $\pm$ 0.21
Claude w.o. RAG ScratchPAD 26.19 $\pm$ 1.31 - 26.09 $\pm$ 0.17 -
Naive RAG 23.46 $\pm$ 0.85  (+2.73) 0.07 $\pm$ 0.00 24.93 $\pm$ 0.20  (+1.16) -4.60 $\pm$ 0.00
APP 22.75 $\pm$ 0.96  (+3.44) 1.53 $\pm$ 0.02 28.16 $\pm$ 0.17  (-2.07) -4.69 $\pm$ 0.01
Gemini w.o. RAG ScratchPAD 25.39 $\pm$ 0.41 - 20.82 $\pm$ 0.01 -
Naive RAG 22.18 $\pm$ 0.39  (+3.21) 0.07 $\pm$ 0.00 23.25 $\pm$ 0.29  (-2.43) -4.60 $\pm$ 0.00
APP 19.78 $\pm$ 0.24  (+5.61) 1.66 $\pm$ 0.09 26.07 $\pm$ 0.05  (-5.24) -4.95 $\pm$ 0.04
Table 2: Validation of CIL estimation. Retrieval number $N=10$. We report mean and std values on twice runs.
We leave the proof in the Appendix A.1. The remaining things are to compute $P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots)$ and $P({\mathcal{X}}_{n_{1}},\cdots)$. Enlightened by Bynum and Cho (2024), we use LLMs to calculate the probabilities. For $P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots)$, note that all ${\mathcal{X}}_{n_{1}}$ have two possible values, namely 0 or 1. We need to sum over all the permutations. We take $P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}}=1,{\mathcal{X}}_{N-2}=0)$ as an example, and derive the prompt from Halawi et al. (2024). We show the prompts in the Appendix A.7 (e). Similar to $P({\mathcal{X}}_{n_{1}},\cdots)$, we take $P({\mathcal{X}}_{n_{1}}=1,{\mathcal{X}}_{n_{2}}=0,)$ for example and use the prompt in the Appendix A.7 (f) to compute. We use window size $w=3$.
Note that LLMs cannot be used to calculate the intervened probability directly since they are trained to be a world model with observation probability Bynum and Cho (2024). We now finish calculating CIL each news article by Equation 3.
4.3 Construction
@@
After calculating the CIL for all pieces of news, we construct the benchmark with them. For each question, we count the number of pieces of news where their CIL are above a threshold. If the number is large enough, we add the question to the chosen set L1, otherwise to L2. We consider L1 to be the main part of our benchmark because answering the questions can be sufficient supported by L1. It can serve as an RAG benchmark. While L2 lacks sufficient support to answer the questions, it also provides valuable information for prediction questions, but needs to be supplemented with additional information beyond the news. We currently create 99 questions for L1 and 53 for L2. We make several discussions about our benchmark:
Data volume. There is not a large volume of valuable prediction questions in total. To ensure the validity of PROPHET, we apply filtering operations during construction by CIL estimation. As a result, the volume of PROPHET is smaller than that of datasets where data collection without inferability validation. This is also the case for other future forecasting datasets with question filtering Karger et al. (2024). We’ll address this issue by using automatic pipelines to regularly collect and add new questions to update the benchmark.
@@
4.4 Statistics and Properties of PROPHET

### v2

@@ 3.2. Causal Intervened Likelihood
For any two news events, the one that occurs later cannot causally affect the one that occurred earlier:
(4) $\forall i,j,\quad\text{if }T_{i}\prec T_{j},\text{then}(X_{j},X_{i})\notin\text{edges~of~SCM}.$
This common-sense assumption aligns causal relationships with the flow of time and eliminates cycles in the SCM. Note that $Y$ is the chronologically last variable in this model.
Assumption 2. $w$-day Dependency Window
The direct causal influence between news events is time-limited. A news event $X_{j}$ can only have a direct causal effect on a news event $X_{i}$ if $X_{j}$ occurs within a $w$-day window before $X_{i}$:
(5) $\forall i,j,\quad\text{if }(T_{i}-T_{j})>w\text{ days, then}(X_{i},X_{j})\notin\text{edges~of~SCM}.$
This assumption posits that the causal influence of distant past news events is mediated by more recent, intermediate news events.
With these assumptions, we can derive the CIL estimation. We first show the calculation for $P(Y=\hat{Y}|\mathrm{do}(X_{i}=1))$; the term $P(Y=\hat{Y}|\mathrm{do}(X_{i}=0))$ can be computed analogously.
@@
The intervened probability $P(Y=\hat{Y}\mid\mathrm{do}(X_{i}=1))$ can be expressed purely in terms of observational probabilities:
(6) $P\big(Y=\hat{Y}\mid\mathrm{do}(X_{i}=1)\big)$ $=\sum_{\mathbf{x}_{\mathcal{N}_{i}}}P\!\left(Y=\hat{Y}\mid X_{i}=1,\ \mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}}\right)$
$\quad\times P\!\left(\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}}\right).$
  • $\mathcal{N}_{i}=\{\,j\mid 0\leq T_{i}-T_{j}\leq w\,\}$ is the set of indices to news articles published within the $w$-day window prior to $X_{i}$.
  • $\mathbf{X}_{\mathcal{N}_{i}}$ denotes the vector of binary random variables $(X_{j})_{j\in\mathcal{N}_{i}}$, where $X_{j}\in\{0,1\}$ indicates the occurrence of $j$th news event.
  • $\mathbf{x}_{\mathcal{N}_{i}}$ denotes one specific assignment of values to $\mathbf{X}_{\mathcal{N}_{i}}$, i.e., a binary vector $(x_{j})_{j\in\mathcal{N}_{i}}$ with each $x_{j}\in\{0,1\}$.
  • The summation $\sum_{\mathbf{x}_{\mathcal{N}_{i}}}$ runs over all $2^{|\mathcal{N}_{i}|}$ possible binary value combinations of $\mathbf{x}_{\mathcal{N}_{i}}$, thereby enumerating every possible pattern of possibility of events in $\mathcal{N}_{i}$.
Proof.
Let $\mathcal{N}_{i}=\{j\mid 0\leq T_{i}-T_{j}\leq w\}$ be the set of indices for variables within the $w$-day window before $X_{i}$, and let $\mathcal{M}_{i}=\{j\mid T_{i}-T_{j}>w\}$ be the set for those outside the window. Using the law of total probability, we can expand the post-intervention probability by marginalizing over all variables that are not descendants of $X_{i}$. Due to Assumption 1 (Temporality), we only need to consider variables that precede $X_{i}$.
(7) $P(Y=\hat{Y}|\mathrm{do}(X_{i}=1))$
$=\sum_{\mathbf{x}_{\mathcal{N}_{i}}}\sum_{\mathbf{x}_{\mathcal{M}_{i}}}P(Y=\hat{Y}|X_{i}=1,\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}},\mathbf{X}_{\mathcal{M}_{i}}=\mathbf{x}_{\mathcal{M}_{i}})$
$\quad\times P(\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}},\mathbf{X}_{\mathcal{M}_{i}}=\mathbf{x}_{\mathcal{M}_{i}}|\mathrm{do}(X_{i}=1)).$
Based on the adjustment formula in causal inference (Pearl, 2010), an intervention on $X_{i}$ only affects its descendants. Due to Assumption 1, the variables $\mathbf{X}_{\mathcal{N}_{i}}$ and $\mathbf{X}_{\mathcal{M}_{i}}$ all occur before $X_{i}$ and thus cannot be its descendants. Therefore, the intervention on $X_{i}$ does not affect their joint probability: $P(\mathbf{X}_{\mathcal{N}_{i}},\mathbf{X}_{\mathcal{M}_{i}}|\mathrm{do}(X_{i}=1))=P(\mathbf{X}_{\mathcal{N}_{i}},\mathbf{X}_{\mathcal{M}_{i}})$.
Furthermore, Assumption 2 states there are no direct causal paths from variables in $\mathbf{X}_{\mathcal{M}_{i}}$ to $Y$ that are not mediated by variables closer to $Y$ (including those in $\mathbf{X}_{\mathcal{N}_{i}}$ and $X_{i}$). This implies that given $X_{i}$ and its more immediate predecessors $\mathbf{X}_{\mathcal{N}_{i}}$, $Y$ becomes conditionally independent of the more distant predecessors $\mathbf{X}_{\mathcal{M}_{i}}$.
(8) $\begin{split}&P(Y=\hat{Y}|X_{i}=1,\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}},\mathbf{X}_{\mathcal{M}_{i}}=\mathbf{x}_{\mathcal{M}_{i}})\\ &\qquad=P(Y=\hat{Y}|X_{i}=1,\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}}).\end{split}$
Substituting this back into Equation (7):
(9) $P(Y=\hat{Y}|\mathrm{do}(X_{i}=1))$
$=\sum_{\mathbf{x}_{\mathcal{N}_{i}}}\sum_{\mathbf{x}_{\mathcal{M}_{i}}}P(Y=\hat{Y}|X_{i}=1,\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})P(\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}},\mathbf{X}_{\mathcal{M}_{i}}=\mathbf{x}_{\mathcal{M}_{i}})$
$=\sum_{\mathbf{x}_{\mathcal{N}_{i}}}P(Y=\hat{Y}|X_{i}=1,\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})\sum_{\mathbf{x}_{\mathcal{M}_{i}}}P(\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}},\mathbf{X}_{\mathcal{M}_{i}}=\mathbf{x}_{\mathcal{M}_{i}}).$
The inner summation over all configurations of $\mathbf{X}_{\mathcal{M}_{i}}$ is simply the marginal probability $P(\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})$.
(10) $P(Y=\hat{Y}|\mathrm{do}(X_{i}=1))=\sum_{\mathbf{x}_{\mathcal{N}_{i}}}$ $P(Y=\hat{Y}|X_{i}=1,\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})$
$\times P(\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}}).$
This completes the proof. ∎
The remaining task is to compute the observational probabilities $P(Y=\hat{Y}|X_{i}=1,\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})$ and $P(\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})$. Inspired by Bynum and Cho (2024), we leverage LLMs to estimate these probabilities. Note that since each $X_{j}\in\mathbf{X}_{\mathcal{N}_{i}}$ is a Bernoulli variable, the summation is over all $2^{|\mathcal{N}_{i}|}$ permutations of their values. We construct prompts similar to the approach of Halawi et al. (2024) to query the LLM for these conditional and joint probabilities. The specific prompts used are detailed in Appendix Prompts (e-f). For our experiments, we use a dependency window of $w=30$ days.
Importantly, trained on vast amounts of observational data, LLM excels at providing observational probabilities. However, LLMs are not strong enough to directly compute intervened probabilities (Bynum and Cho, 2024). Our derivation provides the necessary bridge. By computing both terms and substituting them into Equation 3, we can now calculate the CIL scores for each news article.
Figure 3. In-depth statistics on PROPHET. (a) Distribution of article counts per question. (b) Overall distribution of CIL values across all articles. (c) Article density along event progress (date ranging from the earliest to the latest news) for positive (CIL $\geq$ 0) and negative (CIL ¡ 0). (d) Mean CIL values and standard deviations as a function of forecasting progress.
3.3. News Events Composition
In practical computation of CIL  (Eq. (6)), we observe that the number of news articles within the $w$-day dependency window (we use $w=30$ days) can be extremely large. Directly enumerating all $2^{|\mathcal{N}_{i}|}$ value combinations of $\mathbf{X}_{\mathcal{N}_{i}}$ would lead to exponential computational costs, making the naive implementation infeasible. To address this issue, we design a News Events Composition procedure to aggregate and compress related news items into fewer representative nodes before enumeration. This step preserves the key causal information needed for CIL computation while significantly reducing complexity.
The motivation behind this approach is that many news articles within a short period may describe similar or redundant events, differing only in details or reporting styles, and can be generalized to larger-granularity events. Thus, we aggregate temporally and semantically similar articles into concise, representative summaries.
Formally, when computing CIL for $X_{i}$, we proceed as follows:
  • Individual summarization. We first employ the LLM to produce a concise, factual summary for each news article $X_{j}$ in the dataset. These summaries capture the essential event semantics while discarding irrelevant text details, serving as a standardized basis for grouping.
  • Grouping within the dependency window. We identify the set $\mathcal{N}_{i}=\{j\,|\,0\leq T_{i}-T_{j}\leq w\}$, i.e., all news events occurring within $w$ days before $X_{i}$. We then divide these preceding events into groups of at most $10$ articles each, ordered chronologically.
  • Group-level relevance extraction. For each group, we concatenate the summaries of its articles into a single text block. We then query the LLM to extract and condense only the information relevant to $X_{i}$ from this block. This step removes unrelated details and focuses on causally pertinent content.
  • Clustering of extracted summaries. The relevance-extracted summaries from all groups are embedded into a semantic vector space (e.g., via sentence embeddings [^4] https://sbert.net/) and clustered based on semantic similarity. Clustering captures redundancy among different groups, as reports on similar events may appear in multiple sources or days.
  • Final cluster summarization. For each cluster, we again use the LLM to summarize the shared content into a final, compact representation of that event cluster. These final summaries constitute the reduced set of nodes that we enumerate when computing the CIL  for $X_{i}$.
Through this multi-stage compression pipeline, we replace a large, potentially redundant set $\mathcal{N}_{i}$ with a much smaller set of semantically distinct representative events. This significantly decreases the number of permutations over $\mathbf{X}_{\mathcal{N}_{i}}$ in Equation equation 6, thereby reducing the overall computational load, while still retaining the necessary causal context for accurate interventional probability estimation.
3.4. Construction
@@
After calculating the CIL for all pieces of news, we construct the benchmark with them. For each question, if it contains news articles with CIL¿ 0.7, we add the question to the chosen set L1, otherwise to L2. We consider L1 to be the main part of our benchmark because answering the questions can be sufficiently supported by L1. It can serve as an RAG benchmark. While L2 lacks sufficient support to answer the questions, it also provides valuable information for prediction questions, but needs to be supplemented with additional information beyond the news. We use QwQ-32B (Qwen et al., 2025) to compute all observation probabilities and Qwen2.5-32B (Qwen et al., 2025) to do other data processing. Note that:
@@
3.5. Statistics and Properties of PROPHET